OpenStudy (18jonea):

please help A machine with an AMA of 0.650 and an efficiency of 73% produces a force of 9.52 x 10^6 N over a distance of 2.2 x 10^2 m. What is the input distance? 2.87 x 10^9 m 1.46 x 10^7 m 2.09 x 10^9 m 1.96 x 10^2 m

OpenStudy (18jonea):

@phi

OpenStudy (18jonea):

@ShadowLegendX

OpenStudy (phi):

what's an AMA ?

OpenStudy (18jonea):

actual mechanical advantage

OpenStudy (18jonea):

@phi

OpenStudy (phi):

I have never seen a question like this, and I don't understand it. Do you have more background you can explain?

OpenStudy (18jonea):

@johnweldon1993

OpenStudy (18jonea):

i dont have that much info on this

OpenStudy (phi):

If you have notes that show a similar problem with the answer, please post it.

OpenStudy (18jonea):

A machine with an AMA of 0.35 and an efficiency of 62% produces a force of 5,798 N over a distance of 17.3 m. What is the input distance?

OpenStudy (johnweldon1993):

Ehh, instead of typing everything out...since I see you have another problem which is basically the same...use that as a guideline! Where are you getting messed up?

OpenStudy (phi):

well that answers how to do it. First find the work, Force * distance

OpenStudy (18jonea):

it was just one that was provided, and it doesn't make any sense to me

OpenStudy (18jonea):

so what numbers would i use? these? force of 9.52 x 10^6 N over a distance of 2.2 x 10^2 m

OpenStudy (phi):

yes, that gives the work (or energy)

OpenStudy (johnweldon1993):

Got it Well first, we can agree that a machines efficiency is the ratio of work out to work in \[\large Efficiency = \frac{W_{out}}{W_{in}} \times 100\text{%}\] So from that, we know the efficiency...and we can find the output work! \[\large W_{out} = F\times d\] So here, what would be the output work? \[\large W_{out} = 9.52\times 10^6 [N] \times 2.2\times 10^2 [m] = ?\]

OpenStudy (18jonea):

2.0944 × 10^9

OpenStudy (18jonea):

@johnweldon1993

OpenStudy (johnweldon1993):

Looks right to me, and that is in units of Joules [ J ] Now that is out Output work...so lets plug that into the equation we have so far for efficiency \[\large Efficiency = \frac{W_{out}}{W_{in}}\times 100\text{%}\] We can now rearrange this to solve for the unknown "Input work" \[\large W_{in} = \frac{W_{out}}{Efficiency}\times 100\text{%}\] \[\large W_{in} = \frac{2.0944\times 10^9 [J]}{73\text{%} }\times 100\text{%} = ?\]

OpenStudy (johnweldon1993):

If the percents are throwing you off...this is the same as \[\large W_{in} = \frac{2.0944\times 10^9 [J]}{0.73}\]

OpenStudy (18jonea):

2869041095.89 which is So, the final answer is 2.86904 × 10^9

OpenStudy (johnweldon1993):

Right \(\large W_{in} = 2.87\times 10^9\) *Just rounded off a bit to make it look nicer* Now from here....we recall that the AMA is the ratios of output force to input force...or \(\large AMA = \frac{F_{out}}{F_{in}}\) So we can, from that, solve for the Input force since we know both of the other terms \[\large F_{in} = \frac{F_{out}}{AMA} = \frac{9.52\times 10^6 [N]}{0.650} =?\]

OpenStudy (18jonea):

1.46462 × 10^7

OpenStudy (18jonea):

1.46 X 10^7

OpenStudy (johnweldon1993):

Lol glad you rounded, either way correct...so that is out Input force Now the last step is to solve for what we need Remember how Work = Force times distance? And we used that to solve for the Output work? We're gonna do the same thing here for the inputs \[\large W_{in} = F_{in}\times d_{in}\] We are solving for \(\large d_{in}\) so \[\large d_{in} = \frac{W_{in}}{F_{in}} = \frac{2.87\times 10^9}{1.46\times 10^7}=?\]

OpenStudy (18jonea):

1.96 × 10^2

OpenStudy (johnweldon1993):

And we're done!

OpenStudy (18jonea):

thank you so much!!!!!!

OpenStudy (johnweldon1993):

Not a problem!